OpenStudy (mrnebula):

HELP!!! 3(4 – 2 x ) = –2

OpenStudy (mrnebula):

@tHe_FiZiCx99 @joemarket

OpenStudy (abmon98):

12-6x=-2, -6x=-14 x=14/6

OpenStudy (kewlgeek555):

Okay, just to clarify...your equation is...\[3(4-2x)=-2\]right?

OpenStudy (mrnebula):

yuz

OpenStudy (kewlgeek555):

Well the first step in this is to *distribute* 3 into the parenthesis. This means that you will multiply the numbers inside the parenthesis by 3, because you are distributing 3. So...  You should calculate\[3\times4\]and\[3\times-2\]